# Composite likelihood estimation of stationary Gaussian processes

### v1

@@ 6 Conclusion
In this paper, we develop a framework for doing composite likelihood inference of a parametric class of stationary Gaussian processes. We derive the consistency and asymptotic distribution of the maximum composite likelihood estimator (MCLE) under various degrees of persistence.
@@
In our empirical implementation of the fOU process, we discover that the log-variance of the large-cap segment of the cryptocurrency spot exchange rate market exhibits roughness. On the one hand, this agrees with Gatheral, Jaisson, and Rosenbaum (2018) and the universality of the volatility formation process, see Rosenbaum and Zhang (2022). On the other hand, it stands in contrast to the long memory version of the fOU process, proposed as a model for stochastic volatility in Comte and Renault (1998), and the abundant literature on modeling and forecasting realized variance as a long-range dependent variable, e.g. Andersen, Bollerslev, Diebold, and Labys (2003) and Corsi (2009). This “volatility puzzle” is scrutinized further in Shi and Yu (2022), who note that the likelihood function of a discrete-time version of the fOU is bimodal, leading to near-observational equivalence between these two opposing regimes. Li, Phillips, Shi, and Yu (2023) explore the issue in terms of weak identification and the empirical evidence of their identification robust inference is rooting for long memory.
@@
In our opinion, the source of the dispute is that the fOU specification is not flexible enough. It has to control both the short- and long-term persistence in a single parameter—the Hurst exponent. As it cannot accommodate both, it therefore needs to make a stance: roughness or long memory.
@@
Set against this backdrop, a logical step is to divert more attention toward models that decouple the short- and long-run persistence. There are numerous ways one can do that. A standard approach in the SV literature is to superposition independent driving factors (see, e.g., Alizadeh, Brandt, and Diebold, 2002; Barndorff-Nielsen and Shephard, 2002; Christoffersen, Jacobs, and Mimouni, 2010). However, such multi-factor models usually exhibit exponentially decaying acf and, hence, at best approximate roughness and long memory. An intriguing avenue for further research is to superposition two independent fOU processes and check whether the empirical estimates of the Hurst indexes are located on either side of the memory spectrum.
@@
Gneiting and Schlather (2004) proposed a so-called Cauchy class of Gaussian processes. As the fOU, it has three parameters to fit the acf, but—opposite the fOU—it reserves one to fit the short-run persistence and another one to fit the long-run persistence. Bennedsen, Lunde, and Pakkanen (2022) implement this model with MME on a vast cross-sectional set of empirical high-frequency equity data and find that both roughness and long memory are embedded in the log-variance process. In the present paper, we also adopt this strategy to explore the issue based on transaction price data from the cryptocurrency market. Interestingly, our empirical MCLE of the Cauchy class also point toward roughness and long memory being present in the log-spot variance. Thus, our results can help to reconcile the conflicting evidence in the previous literature.
@@
At the moment, our procedure suffers from the fact that spot volatility is latent. The sampling variation inherent in estimators, such as realized variance, impairs the MCLE and causes a potentially severe degree of spurious roughness. Bolko, Christensen, Pakkanen, and Veliyev (2023) account for this in their GMM estimation of the fOU model. In future research, we should attempt to do this as well.
@@ Appendix A Mathematical proofs
If we let $\theta_{0}$ denote the true data-generating value of $\theta$, the information inequality implies that $C(\theta)$ is uniquely maximized at $\theta_{0}$ (e.g., Newey and McFadden, 1994, Lemma 2.2). This requires $\theta_{0}$ is identified from the density function, which holds by Assumption 2.2. Theorem 4.1 and 4.3 in Wooldridge (1994) then establishes the uniform weak law of large numbers for the sequence of $q$-wise likelihood functions. From this, the consistency of $\hat{\theta}_{\text{MCLE}}$ follows. $\blacksquare$
@@
and $\Sigma_{k^{j}}=\text{cov}(Y_{i}^{k^{j}})$, which is independent of $i$ by stationarity. Here, we emphasize that the dependence on $\theta$ is solely through $\Sigma$.
@@
$\begin{split}s_{n}(\theta)&=\frac{\partial}{\partial\theta}cl(\theta;y)\\ &=\left[\sum_{k^{j}\in Q}\sum_{i=1}^{n-k_{q^{j}}^{j}}\left(-\frac{1}{2}\frac{\partial}{\partial\theta_{r}}\log\det\big{(}\Sigma_{k^{j}}(\theta)\big{)}-\frac{1}{2}(y_{i}^{k^{j}})^{\top}\frac{\partial}{\partial\theta_{r}}\Sigma_{k^{j}}^{-1}(\theta)y_{i}^{k^{j}}\right)\right]_{r=1}^{p}\end{split}$ (A.1)
@@
$-\frac{1}{n}\frac{\partial}{\partial\theta^{\top}}s_{n}(\bar{\theta})\overset{\mathbb{P}}{\longrightarrow}H(\theta_{0}),$
@@
$H(\theta_{0})=-\mathbb{E}\left(\sum_{k^{j}\in Q}\frac{\partial^{2}}{\partial\theta^{\top}\partial\theta}\log f_{k^{j}}(Y_{1}^{k^{j}};\theta)\mid_{\theta=\theta_{0}}\right)$
@@
$(y_{i}^{k^{j}})^{\top}\frac{\partial}{\partial\theta_{r}}\Sigma_{k^{j}}^{-1}(\theta)y_{i}^{k^{j}}=\sum_{j_{1},j_{2}\in k^{j}}y_{i+j_{1}}y_{i+j_{2}}\frac{\partial}{\partial\theta_{r}}(\Sigma_{k^{j}}^{-1}(\theta))_{j_{1},j_{2}},$
where $\big{(}\Sigma_{k^{j}}^{-1}(\theta)\big{)}_{j_{1},j_{2}}$ is shorthand notation for the entry in $\Sigma_{k^{j}}^{-1}(\theta)$ corresponding to the covariance between $Y_{j_{1}}$ and $Y_{j_{2}}$.
@@
$-\frac{1}{2}\sum_{j_{1},j_{2}\in k^{j}}Y_{i+j_{1}}Y_{i+j_{2}}\left[\frac{\partial}{\partial\theta_{r}}\big{(}\Sigma_{k^{j}}^{-1}(\theta)\big{)}_{j_{1},j_{2}}\right]_{r=1}^{p}\Bigg{)},$
@@
$\text{var}\left(\frac{1}{\sqrt{n}}\sum_{k^{j}\in Q}\sum_{i=1}^{n-k_{q^{j}}}\sum_{j_{1},j_{2}\in k^{j}}Y_{i+j_{1}}Y_{i+j_{2}}\right),$
@@
$\begin{split}\text{var}&\left(\sum_{k^{j}\in Q}\sum_{i=1}^{n-k_{q^{j}}}\sum_{j_{1},j_{2}\in k^{j}}Y_{i+j_{1}}Y_{i+j_{2}}\right)\\ &=\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{i_{1}=1}^{n-k^{j_{1}}_{q^{j_{1}}}}\sum_{i_{2}=1}^{n-k^{j_{2}}_{q^{j_{2}}}}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}\mathbb{E}[Y_{i_{1}+\iota_{1}}Y_{i_{1}+\iota_{2}}Y_{i_{2}+\iota_{3}}Y_{i_{2}+\iota_{4}}]-\mathbb{E}[Y_{i_{1}+\iota_{1}}Y_{i_{1}+\iota_{2}}]\mathbb{E}[Y_{i_{2}+\iota_{3}}Y_{i_{2}+\iota_{4}}]\\ &=\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{i_{1}=1}^{n-k^{j_{1}}_{q^{j_{1}}}}\sum_{i_{2}=1}^{n-k^{j_{2}}_{q^{j_{2}}}}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}\big{(}\mathbb{E}[Y_{i_{1}+\iota_{1}}Y_{i_{1}+\iota_{2}}]\mathbb{E}[Y_{i_{2}+\iota_{3}}Y_{i_{2}+\iota_{4}}]\\ &+\mathbb{E}[Y_{i_{1}+\iota_{1}}Y_{i_{2}+\iota_{3}}]\mathbb{E}[Y_{i_{1}+\iota_{2}}Y_{i_{2}+\iota_{4}}]+\mathbb{E}[Y_{i_{1}+\iota_{1}}Y_{i_{2}+\iota_{4}}]\mathbb{E}[Y_{i_{1}+\iota_{2}}Y_{i_{2}+\iota_{3}}]\big{)}-\mathbb{E}[Y_{i_{1}+\iota_{1}}Y_{i_{1}+\iota_{2}}]\mathbb{E}[Y_{i_{2}+\iota_{3}}Y_{i_{2}+\iota_{4}}]\\ &=\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{i_{1}=1}^{n-k^{j_{1}}_{q^{j_{1}}}}\sum_{i_{2}=1}^{n-k^{j_{2}}_{q^{j_{2}}}}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}\gamma_{i_{1}+\iota_{1}-(i_{2}+\iota_{3})}\gamma_{i_{1}+\iota_{2}-(i_{2}+\iota_{4})}+\gamma_{i_{1}+\iota_{1}-(i_{2}+\iota_{4})}\gamma_{i_{1}+\iota_{2}-(i_{2}+\iota_{3})}\\ &\sim\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{l=-n}^{n}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}(n-l)\left(\gamma_{l+\iota_{1}-\iota_{3}}\gamma_{l+\iota_{2}-\iota_{4}}+\gamma_{l+\iota_{1}-\iota_{4}}\gamma_{l+\iota_{2}-\iota_{3}}\right).\end{split}$ (A.2)
@@
$V_{0}$ $=\frac{1}{4}\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{l=-\infty}^{\infty}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}\left[\frac{\partial}{\partial\theta_{r}}\big{(}\Sigma_{k^{j_{1}}}^{-1}(\theta)\big{)}_{\iota_{1},\iota_{2}}|_{\theta=\theta_{0}}\right]_{r=1}^{p}\left(\left[\frac{\partial}{\partial\theta_{r}}\big{(}\Sigma_{k^{j_{2}}}^{-1}(\theta)\big{)}_{\iota_{3},\iota_{4}}|_{\theta=\theta_{0}}\right]_{r=1}^{p}\right)^{\top}$
×$\times\left(\gamma_{l+\iota_{1}-\iota_{3}}\gamma_{l+\iota_{2}-\iota_{4}}+\gamma_{l+\iota_{1}-\iota_{4}}\gamma_{l+\iota_{2}-\iota_{3}}\right),$
@@
var $\left(\sum_{k^{j}\in Q}\sum_{i=1}^{n-k_{q^{j}}}\sum_{j_{1},j_{2}\in k^{j}}Y_{i+j_{1}}Y_{i+j_{2}}\right)$
$\sim\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{l=-n}^{n}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}(n-l)\left(\gamma_{l+\iota_{1}-\iota_{3}}\gamma_{l+\iota_{2}-\iota_{4}}+\gamma_{l+\iota_{1}-\iota_{4}}\gamma_{l+\iota_{2}-\iota_{3}}\right)$
@@
×$\times\left(\frac{L_{\infty}(l+\iota_{1}-\iota_{3})}{\sqrt{|l+\iota_{1}-\iota_{3}|}}\frac{L_{\infty}(l+\iota_{2}-\iota_{4})}{\sqrt{|l+\iota_{2}-\iota_{4}|}}+\frac{L_{\infty}(l+\iota_{1}-\iota_{4})}{\sqrt{|l+\iota_{1}-\iota_{4}|}}\frac{L_{\infty}(l+\iota_{2}-\iota 3)}{\sqrt{|l+\iota_{2}-\iota_{3}|}}\right)$
$\sim n\text{var}(Y)+\sum_{l=1}^{n}(n-l)L_{\infty}^{2}(l)l^{-1},$

### v2

@@ 5.1 Analysis of spot variance
5.2 Analysis of trading volume
There is abundant empirical evidence that the trading volume of an asset is closely connected to its return volatility (e.g., Epps and Epps, 1976; Hasbrouck, 1991). At a theoretical foundation, this observation is embedded in the so-called mixture of distribution hypothesis (e.g., Clark, 1973; Tauchen and Pitts, 1983), where trading volume and volatility are subordinated by a “news” process. A positive correlation between them is also a common prediction in market microstructure theory (e.g., Glosten and Milgrom, 1985; Easley and O’Hara, 1992). In contrast to spot volatility, though, trading volume is observed.
In this section, we apply our MCLE procedure to the trading volume (measured in USD) of the five cryptocurrencies with the aim of establishing whether roughness and long memory is also present in this variable. We follow the implementation from above and first remove intraday periodicity in the 15-second spot trading volume, which is nearly identical to that observed in the 15-second squared log-return process, adding further support to their strong relationship. We cumulate trading volume over each non-overlapping two-hour block matching the RV estimation and log-transform the end product. However, we observe a slight uptick in the log-trading volume over time, which we remove by fitting a regression with a deterministic linear trend to the series.
The right-hand side of Table 5 reports the estimated parameter vector of the fOU process and Cauchy class based on the de-trended and de-seasonalized log-trading volume. The results are more or less in tandem with those obtained for log-RV. In particular, there is strong evidence of roughness in log-trading volume with the roughness index being estimated around 0.35-0.35 to 0.40-0.40 for both models. Moreover, the mean reversion parameter in the fOU model is close to the zero lower bound, again indicative of near-unit root behavior. Furthermore, the memory parameter of the Cauchy class is assessed to be small, pointing to long memory, and is, in fact, even lower than for the log-RV, but it remains squarely within the interior of the parameter space.
The analysis of two-hour log-RV sought to balance the “smoothing” and “roughing” effect in the volatility proxy, while dodging microstructure noise with 15-second sampling of the log-price process. We do not encounter this trade-off with the trading volume, so we can sample it at a faster pace, which can help to gauge whether our conclusion about roughness and long memory is an artifact of temporal aggregation. To shed light on this, we sample trading volume starting at the two-hour horizon and moving down to the 15-second window, and we repeat the MCLE procedure at every step. In the last instance, the optimizer is confronted with $5{,}760\times 2{,}278=13{,}121{,}280$ observations, but it nonetheless manages to converge in a few minutes.
In Figure 6, we present the estimated roughness index and memory parameter ($\kappa$ for the fOU process and $\beta$ for the Cauchy class) as a function of the sampling frequency. Several noteworthy observations emerge. First, the parameter estimates of the five included cryptocurrencies exhibit near-identical behavior, consistent with previous “universality” evidence from other asset classes (e.g., Bennedsen, Lunde, and Pakkanen, 2022; Rosenbaum and Zhang, 2024). Second, the estimated roughness index is slightly decreasing, as we sample more often, and it ends at around $\widehat{\alpha}\approx-0.47$ for the fOU and $\widehat{\alpha}\approx-0.40$ for the Cauchy class. This indicates an extreme degree of roughness over short time intervals, echoing the empirical evidence from volatility in Bolko, Christensen, Pakkanen, and Veliyev (2023). Third, the estimated $\kappa$ and $\beta$ are rather robust to the choice of time aggregation.
In summary, our parameter estimates based on the log-trading volume corroborate the previous application to the log-RV measure.
Figure 6: Parameter estimates for log-trading volume.
Panel A: fOU. Panel B: fOU.
Refer to caption Refer to caption
Panel C: Cauchy. Panel D: Cauchy.
Refer to caption Refer to caption
Note. We show the parameter estimates of $\kappa$ (Panel A) and $\alpha$ (Panel B) for the fOU process and $\beta$ (Panel C) and $\alpha$ (Panel D) for the Cauchy class based on log-trading volume. They are plotted as a function of the sampling frequency going from a two-hour down to a 15-second horizon. Note that the $x$-axis is on a logarithmic scale.
@@ 6 Conclusion
In this paper, we develop a framework for doing composite likelihood estimation of a parametric class of stationary Gaussian processes. We derive the consistency and asymptotic distribution of the maximum composite likelihood estimator (MCLE) under various degrees of persistence.
@@
In our empirical implementation of the fOU process, we discover that the log-variance of the large-cap segment of the cryptocurrency spot exchange rate market exhibits roughness. On the one hand, this agrees with Gatheral, Jaisson, and Rosenbaum (2018) and the universality of the volatility formation process, see Rosenbaum and Zhang (2024). On the other hand, it stands in contrast to the long memory version of the fOU process, proposed as a model for stochastic volatility in Comte and Renault (1998), and the abundant literature on modeling and forecasting RV as a long-range dependent variable, e.g. Andersen, Bollerslev, Diebold, and Labys (2003) and Corsi (2009). This “volatility puzzle” is scrutinized further in Shi and Yu (2023), who note that the likelihood function of a discrete-time version of the fOU is bimodal, leading to near-observational equivalence between these two opposing regimes. Li, Phillips, Shi, and Yu (2025) explore the issue in terms of weak identification and the empirical evidence of their identification robust inference is rooting for long memory.
@@
In our opinion, the source of the dispute is that the fOU specification is not flexible enough. It has to control both the short- and long-term persistence in a single parameter—the Hurst exponent. As it cannot accommodate both, it therefore needs to make a stance: either roughness or long memory.
@@
Set against this backdrop, a logical step is to divert more attention toward models that decouple the short- and long-run persistence. There are numerous ways one can do that. A standard approach in the SV literature is to superposition independent driving factors (see, e.g., Alizadeh, Brandt, and Diebold, 2002; Barndorff-Nielsen and Shephard, 2002). However, such multi-factor models usually exhibit exponentially decaying ACF and, hence, at best approximate roughness and long memory. An intriguing avenue for further research is to superposition two independent fOU processes and check whether the empirical estimates of the Hurst indexes are located on either side of the memory spectrum. This idea is being pursued in Christensen, Espinosa, Pakkanen, and Veliyev (2025), where the GMM approach of Bolko, Christensen, Pakkanen, and Veliyev (2023) is extended to a two-factor setting.
@@
Gneiting and Schlather (2004) proposed a so-called Cauchy class of Gaussian processes. As the fOU, it has two parameters to fit the ACF, but—opposite the fOU—it reserves one to fit the short-run persistence and another one to fit the long-run persistence. Bennedsen, Lunde, and Pakkanen (2022) implement this model with MME on a vast cross-sectional set of empirical high-frequency equity data and find that roughness and long memory are embedded in the log-variance process. In the present paper, we also adopt this strategy to explore the issue based on transaction price data from the cryptocurrency market. Interestingly, our empirical MCLE of the Cauchy class also point toward both roughness and long memory being present in the log-spot variance. Thus, our results can help to reconcile the conflicting evidence in the previous literature.
@@
At the moment, our procedure suffers from the fact that spot volatility is latent. The sampling variation inherent in estimators, such as RV, impairs the MCLE and causes a potentially severe degree of spurious roughness. Bolko, Christensen, Pakkanen, and Veliyev (2023) account for this in their GMM estimation of the fOU model. In future research, we should attempt to do this as well. In lieu of this, we here include an application to the log-trading volume, which confirms our findings for the log-variance.
@@ Appendix A Mathematical proofs
Proof of Lemma 2.1.
In the proof, we look at a single tuple, $k$, of length $q$ only. The general case follows by extension of this result to the sum defining the log-composite likelihood. As explained in the proof of Theorem 2.1, a law of large number implies that, uniformly in $\theta$,
$\frac{1}{n}cl(\theta;Y)\overset{\mathbb{P}}{\longrightarrow}\mathbb{E}_{\theta_{0}}\left(\log f_{k}(Y^{k};\theta)\right),$
as $n\rightarrow\infty$, where $Y^{k}$ denotes the selected $k$-tuple of the random vector $Y$.
Now, recall the information inequality:
$\mathbb{E}_{\theta_{0}}\left[\log\frac{f_{k}(Y^{k};\theta)}{f_{k}(Y^{k};\theta_{0})}\right]\leq\log\mathbb{E}_{\theta_{0}}\left[\frac{f_{k}(Y^{k};\theta)}{f_{k}(Y^{k};\theta_{0})}\right]=0,$
by Jensen’s inequality. This implies that
$\mathbb{E}_{\theta_{0}}\left[\log f_{k}(Y^{k};\theta)\right]\leq\mathbb{E}_{\theta_{0}}\left[\log f_{k}(Y^{k};\theta_{0})\right].$
As the natural logarithm is strictly concave, the above inequality is also strict if $\mathbb{P}\big(\log f_{k}(Y^{k};\theta)=\log f_{k}(Y^{k};\theta_{0})\big)\neq 1$. Recall now that
$\log f_{k}(y;\theta)=C-\frac{1}{2}\log|\Sigma(\theta)|-\frac{1}{2}y^{\top}\Sigma^{-1}(\theta)y.$
This expression depends on $\theta$ only through $\gamma_{h}(\theta)$, and viewed as a function of $y$ it is a second-order polynomial with an intercept. As our assumption implies that for $\theta\neq\theta_{0}$ there exists at least one $h$ such that $\gamma_{h}(\theta_{0})\neq\gamma_{h}(\theta)$, it follows that $\Sigma(\theta)\neq\Sigma(\theta_{0})$, and, therefore, $\Sigma(\theta)^{-1}\neq\Sigma(\theta_{0})^{-1}$ by uniqueness of the inverse. The proof is then completed by noting that a quadratic function can agree for at most finitely many (in fact, two) values of $y$. $\blacksquare$
@@
If we let $\theta_{0}$ denote the true data-generating value of $\theta$, the information inequality implies that $C(\theta)$ is uniquely maximized at $\theta_{0}$ (e.g., Newey and McFadden, 1994, Lemma 2.2). This requires $\theta_{0}$ is identified from the density function, which holds by Assumption 2.2. Theorem 4.1 and 4.3 in Wooldridge (1994) then establishes the uniform weak law of large numbers for the sequence of $q$-wise likelihood functions, from which it follows that $\hat{\theta}_{\text{MCLE}}\overset{\mathbb{P}}{\longrightarrow}\theta_{0}$. $\blacksquare$
@@
and $\Sigma_{k^{j}}=\text{cov}(Y_{i}^{k^{j}})$, which is independent of $i$ by stationarity. Here, we again emphasize that the dependence on $\theta$ is solely through $\Sigma(\theta)$.
@@
$\begin{split}s_{n}(\theta)&=\frac{\partial}{\partial\theta}cl(\theta;y)\\ &=\left[\sum_{k^{j}\in Q}\sum_{i=1}^{n-k_{q^{j}}^{j}}\left(-\frac{1}{2}\frac{\partial}{\partial\theta_{r}}\log\det\big(\Sigma_{k^{j}}(\theta)\big)-\frac{1}{2}(y_{i}^{k^{j}})^{\top}\frac{\partial}{\partial\theta_{r}}\Sigma_{k^{j}}^{-1}(\theta)y_{i}^{k^{j}}\right)\right]_{r=1}^{p}.\end{split}$ (A.1)
@@
$\frac{1}{n}\frac{\partial}{\partial\theta^{\top}}s_{n}(\bar{\theta})\overset{\mathbb{P}}{\longrightarrow}H(\theta_{0}),$
@@
$H(\theta_{0})=-\mathbb{E}\left(\sum_{k^{j}\in Q}\frac{\partial^{2}}{\partial\theta\partial\theta^{\top}}\log f_{k^{j}}(Y_{1}^{k^{j}};\theta)\mid_{\theta=\theta_{0}}\right)$
@@
$(Y_{i}^{k^{j}})^{\top}\frac{\partial}{\partial\theta_{r}}\Sigma_{k^{j}}^{-1}(\theta)Y_{i}^{k^{j}}=\sum_{j_{1},j_{2}\in k^{j}}Y_{(i+j_{1})\Delta}Y_{(i+j_{2})\Delta}\frac{\partial}{\partial\theta_{r}}(\Sigma_{k^{j}}^{-1}(\theta))_{j_{1},j_{2}},$
where $\big(\Sigma_{k^{j}}^{-1}(\theta)\big)_{j_{1},j_{2}}$ is shorthand notation for the entry in $\Sigma_{k^{j}}^{-1}(\theta)$ corresponding to the covariance between $Y_{j_{1}\Delta}$ and $Y_{j_{2}\Delta}$.
@@
$-\frac{1}{2}\sum_{j_{1},j_{2}\in k^{j}}Y_{(i+j_{1})\Delta}Y_{(i+j_{2})\Delta}\left[\frac{\partial}{\partial\theta_{r}}\big(\Sigma_{k^{j}}^{-1}(\theta)\big)_{j_{1},j_{2}}\right]_{r=1}^{p}\Bigg),$
@@
$\text{var}\left(\frac{1}{\sqrt{n}}\sum_{k^{j}\in Q}\sum_{i=1}^{n-k_{q^{j}}}\sum_{j_{1},j_{2}\in k^{j}}Y_{(i+j_{1})\Delta}Y_{(i+j_{2})\Delta}\right),$
@@
$\begin{split}\text{var}&\left(\sum_{k^{j}\in Q}\sum_{i=1}^{n-k_{q^{j}}}\sum_{j_{1},j_{2}\in k^{j}}Y_{(i+j_{1})\Delta}Y_{(i+j_{2})\Delta}\right)\\ &=\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{i_{1}=1}^{n-k^{j_{1}}_{q^{j_{1}}}}\sum_{i_{2}=1}^{n-k^{j_{2}}_{q^{j_{2}}}}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}\mathbb{E}[Y_{(i_{1}+\iota_{1})\Delta}Y_{(i_{1}+\iota_{2})\Delta}Y_{(i_{2}+\iota_{3})\Delta}Y_{(i_{2}+\iota_{4})\Delta}]\\ &-\mathbb{E}[Y_{(i_{1}+\iota_{1})\Delta}Y_{(i_{1}+\iota_{2})\Delta}]\mathbb{E}[Y_{(i_{2}+\iota_{3})\Delta}Y_{(i_{2}+\iota_{4})\Delta}]\\ &=\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{i_{1}=1}^{n-k^{j_{1}}_{q^{j_{1}}}}\sum_{i_{2}=1}^{n-k^{j_{2}}_{q^{j_{2}}}}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}\big(\mathbb{E}[Y_{(i_{1}+\iota_{1})\Delta}Y_{(i_{1}+\iota_{2})\Delta}]\mathbb{E}[Y_{(i_{2}+\iota_{3})\Delta}Y_{(i_{2}+\iota_{4})\Delta}]\\ &+\mathbb{E}[Y_{(i_{1}+\iota_{1})\Delta}Y_{(i_{2}+\iota_{3})\Delta}]\mathbb{E}[Y_{(i_{1}+\iota_{2})\Delta}Y_{(i_{2}+\iota_{4})\Delta}]+\mathbb{E}[Y_{(i_{1}+\iota_{1})\Delta}Y_{(i_{2}+\iota_{4})\Delta}]\mathbb{E}[Y_{(i_{1}+\iota_{2})\Delta}Y_{(i_{2}+\iota_{3})\Delta}]\big)\\ &-\mathbb{E}[Y_{(i_{1}+\iota_{1})\Delta}Y_{(i_{1}+\iota_{2})\Delta}]\mathbb{E}[Y_{(i_{2}+\iota_{3})\Delta}Y_{(i_{2}+\iota_{4})\Delta}]\\ &=\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{i_{1}=1}^{n-k^{j_{1}}_{q^{j_{1}}}}\sum_{i_{2}=1}^{n-k^{j_{2}}_{q^{j_{2}}}}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}\gamma_{(i_{1}+\iota_{1}-(i_{2}+\iota_{3}))\Delta}\gamma_{(i_{1}+\iota_{2}-(i_{2}+\iota_{4}))\Delta}\\ &+\gamma_{(i_{1}+\iota_{1}-(i_{2}+\iota_{4}))\Delta}\gamma_{(i_{1}+\iota_{2}-(i_{2}+\iota_{3}))\Delta}\\ &\sim\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{l=-n}^{n}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}(n-l)\left(\gamma_{(l+\iota_{1}-\iota_{3})\Delta}\gamma_{(l+\iota_{2}-\iota_{4})\Delta}+\gamma_{(l+\iota_{1}-\iota_{4})\Delta}\gamma_{)l+\iota_{2}-\iota_{3})\Delta}\right),\end{split}$ (A.2)
@@
$V_{0}$ $=\frac{1}{4}\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{l=-\infty}^{\infty}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}\left[\frac{\partial}{\partial\theta_{r}}\big(\Sigma_{k^{j_{1}}}^{-1}(\theta)\big)_{\iota_{1},\iota_{2}}|_{\theta=\theta_{0}}\right]_{r=1}^{p}$
$\times\left(\left[\frac{\partial}{\partial\theta_{r}}\big(\Sigma_{k^{j_{2}}}^{-1}(\theta)\big)_{\iota_{3},\iota_{4}}|_{\theta=\theta_{0}}\right]_{r=1}^{p}\right)^{\top}\left(\gamma_{(l+\iota_{1}-\iota_{3})\Delta}\gamma_{(l+\iota_{2}-\iota_{4})\Delta}+\gamma_{(l+\iota_{1}-\iota_{4})\Delta}\gamma_{(l+\iota_{2}-\iota_{3})\Delta}\right),$
@@
var $\left(\sum_{k^{j}\in Q}\sum_{i=1}^{n-k_{q^{j}}}\sum_{j_{1},j_{2}\in k^{j}}Y_{(i+j_{1})\Delta}Y_{(i+j_{2})\Delta)\Delta}\right)$
$\sim\sum_{k^{j_{1}},k^{j_{2}}\in Q}\sum_{l=-n}^{n}\sum_{\iota_{1},\iota_{2}\in k^{j_{1}}}\sum_{\iota_{3},\iota_{4}\in k^{j_{2}}}(n-l)\left(\gamma_{(l+\iota_{1}-\iota_{3})\Delta}\gamma_{(l+\iota_{2}-\iota_{4})\Delta}+\gamma_{(l+\iota_{1}-\iota_{4})\Delta}\gamma_{(l+\iota_{2}-\iota_{3})\Delta}\right)$
@@
$\times\left(\frac{L_{\infty}((l+\iota_{1}-\iota_{3})\Delta)}{\sqrt{|l+\iota_{1}-\iota_{3}|\Delta}}\frac{L_{\infty}((l+\iota_{2}-\iota_{4})\Delta)}{\sqrt{|l+\iota_{2}-\iota_{4}|\Delta}}+\frac{L_{\infty}((l+\iota_{1}-\iota_{4})\Delta)}{\sqrt{|l+\iota_{1}-\iota_{4}|\Delta}}\frac{L_{\infty}((l+\iota_{2}-\iota_{3})\Delta)}{\sqrt{|l+\iota_{2}-\iota_{3}|\Delta}}\right)$
$\sim n\text{var}(Y)+\sum_{l=1}^{n}(n-l)L_{\infty}^{2}(l\Delta)(l\Delta)^{-1},$